Find the pressure exerted on a person by a few plastic blocks of total mass 42 kg, the bottom of which in contact with the person has an area of 840 cm2.

1
Expert's answer
2022-04-19T12:05:58-0400
"P=\frac{F}{A}=\frac{mg}{A}"

"P=\frac{42*9.8}{840*10^{-4}}=4900\:\rm Pa"